OpenStudy (anonymous):

y'=? , y=x.(root of 8-x^2)

OpenStudy (anonymous):

\[y=x \sqrt{8-x ^{2}} \] what is y' and y''

OpenStudy (anonymous):

use the product rule,\[\frac{ d(g(x).f(x)) }{ dx }=g'(x)f(x)+g(x)f'(x)\]\[\times f(x)\]

OpenStudy (anonymous):

ignore "x f(x)" at the bottom.

OpenStudy (anonymous):

huh hahaha okey

OpenStudy (anonymous):

what about the root

OpenStudy (anonymous):

Use,\[ \frac{ d((f(x))^n) }{ dx}=n.f'(x).(f(x))^{ n-1 }\]

OpenStudy (anonymous):

i stoped at \[\sqrt{8-x2}+x \frac{ -2x }{ 2\sqrt{8-x ^{2}} }=0\]

OpenStudy (anonymous):

correct except it doesnt nessecarily =0, it is y'=

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

solved it x=0 ,thanks

OpenStudy (anonymous):

Have you managed y''?